OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 9 where f of x equals x plus 9 when x is less than 9 and f of x equals 9 minus x when x is greater than or equal to 9

OpenStudy (anonymous):

I don't quite understand how they want me to solve this

hartnn (hartnn):

The limit will exist is its value when x approaches from either side, less than 9 or greater than 9, is SAME! so can you tell me whats the limits value when x<9 ?

OpenStudy (anonymous):

I honestly don't know

OpenStudy (anonymous):

is it everything less than 9?

hartnn (hartnn):

do you know how to evaluate limits? like this one? \(\lim \limits_{x\to 9}x+9 =.. ?\) you know what to do?

OpenStudy (anonymous):

yah

hartnn (hartnn):

what ?

OpenStudy (anonymous):

its 18 right?

OpenStudy (anonymous):

the limit to x+9 is 18

OpenStudy (anonymous):

as the values of x approach nine

hartnn (hartnn):

yes! that was when x<9 so you just plug in x = 9 whenever you can. similarly whats \( \lim \limits_{x \to 9 } 9-x = .. ?\)

OpenStudy (anonymous):

that would be zero?

hartnn (hartnn):

correct, now are these same?

OpenStudy (anonymous):

you are plugging in 9 into the x values to get closer to zero?

hartnn (hartnn):

we are plugging in x= 9 in the function, because x approaches 9 when x<9, we got the limit as 18, when x is > or = 9, we got the limit as 0 its like a sudden spike!